OpenStudy (anonymous):

Evaluate the following expression using the values given: 3x - y - 3z if x = -2, y = 1, and z = -2.

OpenStudy (anonymous):

Plug in the numbers: 3 x -2 - 1 - 3 x -2 Use your calculator to multiply all of them: So your answer is -1.

OpenStudy (aayushi.somani):

place the no. in place of x, y and z.. then, 3x-y-3z = 3(-2)-1-3(-2) =-6-1+6 =-1